# Vertex-minor universality of a random graph

### 13 Mar 2026 v1

@@ Abstract.
Given a graph $G$ and a vertex $v\in V(G)$, a local complementation at $v$ on $G$ is an operation that replaces the induced graph on the neighborhood of $v$ by its complement. A graph $H$ is a vertex-minor if $H$ can be obtained from $G$ by a sequence of vertex deletions and local complementation. A graph is said to be $k$-vertex-minor universal if it contains every $k$-vertex graph on any $k$-subset of vertices as a vertex minor. Previously, Ascoli–Fredrickson–Fredrickson–McFarland–Post [undef] proved that with high probability $G(n,1/2)$ is $\Omega(\sqrt{n})$-vertex-minor universal. Furthermore, they conjectured that with high probability $G(n,p)$ and G(n,1p)G(n,1-p) are $\Omega(p\sqrt{n})$-vertex-minor universal for all $\omega(1/\sqrt{n})\leqslant p\leqslant 1/2$. In this short note, we confirm this conjecture up to an extra logarithm factor and show that this is true with probability $1-2^{-\Omega(p^{2}n)}$ if $\Omega(\log n/\sqrt{n})\leqslant p\leqslant 1/2$.
@@ 1. Introduction
Motivated by the connections between vertex-minors and quantum communication, a recent line of work considers the universality question for vertex-minors. A graph $G$ is $k$-vertex-minor universal if for every vertex subset $U\subseteq V(G)$ of size $k$, every graph on $U$ is a vertex-minor of $G$. Note that this is a strong condition where every graph on $k$ vertices must be realized on any specified subset of $k$ vertices in $G$. In this direction, Claudet, Mhalla, and Perdrix [undefc] first showed that there exists $k$-vertex-minor universal graphs with $O(k^{4}\log k)$ vertices and any $k$-vertex-minor universal graph must contain at least $(k-2)^{2}/(2\log_{2}3)$ vertices. Then, Cautrés, Claudet, Mhalla, Perdrix, Savin, and Thomassé [undefb] further showed the existence of $k$-vertex-minor universal graphs with $(2+o(1))k^{2}$ vertices. Most recently, Ascoli, Fredrickson, Fredrickson, McFarland, and Post [undef] showed that almost all graphs on $n$ vertices are $O(\sqrt{n})$-vertex-minor universal by proving the following statement. Recall that $G(n,p)$ denotes the Erdős-Renyí random graph on $n$ vertices with each edge present with probability $p$.
@@ 4. Concluding remarks
where $E^{\prime}$ is the edges of the unordered bipartite graph $(L,R,E)\times uv$. For any ordered bipartite graph $H=(V_{L},V_{R},E(H))$ with $U_{L}\cup U_{R}\subseteq V(G)$, we say that $H$ is a (labeled) pivot-minor of $G$ if it can be obtained on $V_{L}\cup V_{R}$ via a sequence of vertex deletions and pivots. We say that ordered bipartite graph $G$ is $k$-pivot-minor universal if any ordered bipartite graph $H$ on any $k$ vertices of $V(G)$ is a pivot minor of $G$.
Let $G(\ell,r,p)$ denote the random ordered bipartite graph on vertex bipartition $L\sqcup R$ with $|L|=\ell$, $|R|=r$, and edges between $L\times R$ present independently with probability $p$. We believe Lemma 3.1 can be adapted in this setting with the following strategy.

### 20 Mar 2026 v2

@@ Abstract.
Given a graph $G$ and a vertex $v\in V(G)$, a local complementation at $v$ on $G$ is an operation that replaces the induced graph on the neighborhood of $v$ by its complement. A graph $H$ is a vertex-minor if $H$ can be obtained from $G$ by a sequence of vertex deletions and local complementation. A graph is said to be $k$-vertex-minor universal if it contains every $k$-vertex graph on any $k$-subset of vertices as a vertex minor. Previously, Ascoli–Frederickson–Frederickson–McFarland–Post [undef] proved that with high probability $G(n,1/2)$ is $\Omega(\sqrt{n})$-vertex-minor universal. Furthermore, they conjectured that with high probability $G(n,p)$ and G(n,1p)G(n,1-p) are $\Omega(p\sqrt{n})$-vertex-minor universal for all $\omega(1/\sqrt{n})\leqslant p\leqslant 1/2$. In this short note, we confirm this conjecture up to an extra logarithm factor and show that this is true with probability $1-2^{-\Omega(p^{2}n)}$ if $\Omega(\log n/\sqrt{n})\leqslant p\leqslant 1/2$.
@@ 1. Introduction
Motivated by the connections between vertex-minors and quantum communication, a recent line of work considers the universality question for vertex-minors. A graph $G$ is $k$-vertex-minor universal if for every vertex subset $U\subseteq V(G)$ of size $k$, every graph on $U$ is a vertex-minor of $G$. Note that this is a strong condition where every graph on $k$ vertices must be realized on any specified subset of $k$ vertices in $G$. In this direction, Claudet, Mhalla, and Perdrix [undefc] first showed that there exists $k$-vertex-minor universal graphs with $O(k^{4}\log k)$ vertices and any $k$-vertex-minor universal graph must contain at least $(k-2)^{2}/(2\log_{2}3)$ vertices. Then, Cautrés, Claudet, Mhalla, Perdrix, Savin, and Thomassé [undefb] further showed the existence of $k$-vertex-minor universal graphs with $(2+o(1))k^{2}$ vertices. Most recently, Ascoli, Frederickson, Frederickson, McFarland, and Post [undef] showed that almost all graphs on $n$ vertices are $O(\sqrt{n})$-vertex-minor universal by proving the following statement. Recall that $G(n,p)$ denotes the Erdős-Renyí random graph on $n$ vertices with each edge present with probability $p$.
@@ 4. Concluding remarks
where $E^{\prime}$ is the edges of the unordered bipartite graph $(L,R,E)\times uv$. For any ordered bipartite graph $H=(V_{L},V_{R},E(H))$ with $V_{L}\cup V_{R}\subseteq V(G)$, we say that $H$ is a (labeled) pivot-minor of $G$ if it can be obtained on $V_{L}\cup V_{R}$ via a sequence of vertex deletions and pivots. We say that ordered bipartite graph $G$ is $k$-pivot-minor universal if any ordered bipartite graph $H$ on any $k$ vertices of $V(G)$ is a pivot minor of $G$.
Let $G(a,b,p)$ denote the random ordered bipartite graph on vertex bipartition $L\sqcup R$ with $|L|=a$, $|R|=b$, and edges between $L\times R$ present independently with probability $p$. We believe Lemma 3.1 can be adapted in this setting with the following strategy.
@@
After we put the first version of this note on arXiv, the following solution to the question is pointed out to us by Mathieu Rundström. The answer is to this question is affirmative, and it is true that $r^{\prime}\geqslant r/2$ with probability at least $1-2^{-\Omega(qr)}$, where $q=\min(p,1-p)$. This can be proved with the following observation. Let $A$ be the biadjacency matrix of $G$, then the maximum possible $r^{\prime}$ is at least $\gamma\overset{\rm{def}}{=}\textup{rank}(A)$. Here, we recall that $\textup{rank}(A)$ denotes the rank over $\mathbb{F}_{2}$. To see this observation, we can restrict $A$ to an invertible $\gamma\times\gamma$ submatrix $B$. Let $V_{B}$ be the vertex set corresponding to the rows and columns of $B$. It is sufficient to pick $w_{1}w^{\prime}_{1},\dots,w_{\gamma}w^{\prime}_{\gamma}$ from $V_{B}$.
We can show this by inducting on $\gamma$. We first pick an arbitrary entry with value one, say $B_{11}=1$, and set $w_{1}w^{\prime}_{1}$ to be the edge correspond to this entry. Suppose $B=\begin{pmatrix}1&\alpha\\ \beta&B^{\prime}\end{pmatrix}$, where $\alpha\in\mathbb{F}_{2}^{1\times(\gamma-1)}$, $\beta\in\mathbb{F}_{2}^{(\gamma-1)\times 1}$, and $B^{\prime}\in\mathbb{F}_{2}^{(\gamma-1)\times(\gamma-1)}$. It follows that the biadjacency matrix of $(G\times w_{1}w^{\prime}_{1})[V_{B}]$ equals $\begin{pmatrix}1&0\\ 0&B^{\prime}+\alpha\beta\end{pmatrix}$. Note that $B^{\prime}+\alpha\beta$ is invertible since
$\begin{pmatrix}1&0\\ 0&B^{\prime}+\alpha\beta\end{pmatrix}=\begin{pmatrix}1&0\\ \beta&I_{\gamma-1}\end{pmatrix}\begin{pmatrix}1&\alpha\\ \beta&B^{\prime}\end{pmatrix}\begin{pmatrix}1&\alpha\\ 0&I_{\gamma-1}\end{pmatrix}$
is invertible. Thus, we can find $w_{2}w^{\prime}_{2},\dots,w_{\gamma}w^{\prime}_{\gamma}$ by applying the inductive hypothesis to $B^{\prime}+\alpha\beta$.
Now, it is sufficient to show that with high probability $\textup{rank}(A)\geqslant r/2$. Note that $A\in\mathbb{F}_{2}^{r\times r}$ is a random binary matrix with i.i.d. $\textup{Ber}(p)$ entries. Let $A_{i}$ be the $i$-th column vector of $A$. We first bound the probability $\mathbb{P}[A_{i+1}\in\textup{span}(A_{1},\dots,A_{i})\mid A_{1},\dots,A_{i}]$. Let $\mathcal{V}_{i}=\textup{span}(A_{1},\dots,A_{i})$ and suppose the dimension of $\mathcal{V}_{i}$ is $d_{i}$. Given $A_{1},\dots,A_{i}$, we can find a set $I_{i}$ of $d_{i}$ entries such that the projection $\pi_{I_{i}}(\mathcal{V}_{i})$ onto $I_{i}$ is surjective. Hence, for any $v\in\mathcal{V}_{i}$, the projection $\pi_{[r]\setminus I_{i}}(v)$ is uniquely determined by $\pi_{I_{i}}(v)$. Indeed, we know that there is a linear map $T$, depending on $\mathcal{V}_{i}$, such that $\pi_{[r]\setminus I_{i}}(v)=T(\pi_{I_{i}}(v))$. Thus, $A_{i+1}\in\mathcal{V}_{i}$ only if $\pi_{[r]\setminus I_{i}}(A_{i+1})=T(\pi_{I_{i}}(A_{i+1}))$, which happens with probability at most $(1-q)^{r-d_{i}}$ since we can first expose $\pi_{I_{i}}(A_{i+1})$. Hence, we can conclude that
$\mathbb{P}[A_{i+1}\in\mathcal{V}_{i}\mid A_{1},\dots,A_{i}]\leqslant(1-q)^{r-d_{i}}.$
Fix $0<\gamma_{0}<r$. Let $\mathbbm{1}_{i}$ be the indicator function of the event that $A_{i}\in\mathcal{V}_{i-1}$ and $\dim\mathcal{V}_{i-1}\leqslant\gamma_{0}$ for $i=1,\dots,r$, assuming that $\mathcal{V}_{0}=\{0\}$. Set $Z=\sum_{i=1}^{r}\mathbbm{1}_{i}$ to be the number of indices $i$ such that $A_{i}\in\mathcal{V}_{i-1}$ happens before the rank exceeds $\gamma_{0}$. We know that $\mathbb{E}[\mathbbm{1}_{i}]\leqslant(1-q)^{r-d_{i}}\leqslant(1-q)^{r-\gamma_{0}}$ holds for all $i=1,\dots,r$. Thus, by linearity of expectation we have $\mathbb{E}[Z]\leqslant r(1-q)^{r-\gamma_{0}}$. On the other hand, we know that $\textup{rank}(A)\leqslant\gamma_{0}$ only if $Z\geqslant r-\gamma_{0}$. Thus,
$\mathbb{P}[\textup{rank}(A)\leqslant\gamma_{0}]\leqslant\mathbb{P}[Z\geqslant r-\gamma_{0}]\leqslant\frac{\mathbb{E}[Z]}{r-\gamma_{0}}\leqslant\frac{r(1-q)^{r-\gamma_{0}}}{r-\gamma_{0}}.$
If we set $\gamma_{0}=r/2$, then the probability is at most $2(1-q)^{r/2}=2^{-\Omega(qr)}$. This proves the claim that $r^{\prime}\geqslant r/2$ with probability at least $1-2^{-\Omega(qr)}$.
From all the discussion above, we can conclude in a way similar to Section 3 that the bipartite graph $G$ is pivot-minor universal with high probability when $\left|L\right|,\left|R\right|$ are not too unbalanced. To be more explicit, suppose $\left|L\right|=a,\left|R\right|=b$ with $a\geqslant b$, and $G\sim G(L,R,p)$. We can show that $G$ is $k$-pivot-minor universal with high probability if $k=O(\min(q\sqrt{b},q^{2}b/\log a))$. This can be done by picking $s=\Theta(q^{2}b)$ and $r=\left|R\right|-s=\Theta(b)$ in the discussion above. We know that for any fixed ordered bipartite graph $H=(V_{L},V_{R},E(H))$ with $V_{L}\cup V_{R}\subseteq V(G)$ and $\left|V_{L}\cup V_{R}\right|=k$, the probability that $H$ is not a pivot minor of $G$ is $O(2^{-q^{2}b})$ since $k=O(q\sqrt{b})$. Because there are at most $2^{O(k\log a)}$ many choices for $(V_{L},V_{R})$, we can conclude by the union bound that the probability that $G$ is not $k$-pivot-minor universal is at most $2^{O(k\log a)-\Omega(q^{2}b)}=o(1)$.